Find the slope and y-intercept of the following line. Write fractional answers in lowest terms. 5y+3x+2=19+3x.
Answer by ewatrrr(24785)


Hi
Finding the slope and y-intercept of the following line.
5y+3x+2=19+3x
5y = 17
y = 17/5 |slope m = 0 and y-intercept is Pt(0,17/5)